# Weak measurement-based protocol for ergotropy protection in open quantum batteries

### v1

@@ V Acknowledgements
The authors thank Michał Horodecki and Paweł Mazurek for the valuable discussions and suggestions. A.H.A.M. and R.S. acknowledge support from National Science Centre, Poland Grant OPUS-21 (No. 2021/41/B/ST2/03207). B.A. acknowledges support from National Science Center, Poland within the QuantERA II Programme (No 2021/03/Y/ST2/00178, acronym ExTRaQT) that has received funding from the European Union’s Horizon 2020. P.R.D. acknowledges support from the NCN Poland, ChistEra-2023/05/Y/ST2/00005 under the project Modern Device Independent Cryptography (MoDIC).
@@ Appendix A Dissipation and battery discharging
This phenomenon is illustrated in Fig. 10 for various coherent and incoherent initial states, characterized by $P_{0}$ and $Q_{0}$. Notably, when the initial state is active, the ergotropy decreases monotonically and irreversibly over time. In contrast, for passive states, the ergotropy remains constant and zero, as indicated by the red dashed line. Additionally, note that the discharging rates vary significantly. For instance, in the case of a fully charged battery, represented by the solid black line, the decay is faster than for initially coherent states. From this perspective, certain partially charged initial states may retain more ergotropy after a period of dissipation, presenting a potential advantage. This counterintuitive behavior arises from the complex interplay between the incoherent and coherent components of ergotropy during thermalization.
Figure 10: Discharging process: Ergotropy behavior as a function of the dissipation time for distinct initial states, characterized by $P_{0}$ and $Q_{0}$. It was assumed $\omega=1$, $f=0.3$ and $\gamma=10^{-2}$.
@@
Figure 11: Incoherent ergotropy in terms of the dissipation time and initial population $P_{0}$. The dashed red line highlights the curve $\tau_{1/2}$, representing the instant where the diagonal state becomes passive ($P(\tau_{1/2})=1/2$). It was assumed $\omega=1$, $f=0.3$, $\gamma=10^{-2}$.
On the other hand the coherent ergotropy at time $t$ is written as [97, 98]
$\mathcal{R}^{coh}[\hat{\rho}(t)]=\frac{1}{2}\left(\psi(t)-\sqrt{\psi^{2}(t)-4\left|Q(t)\right|^{2}}\right),$ (44)
where $\psi(t)=\sqrt{2\mu[\hat{\rho}(t)]-1}$, with $\mu[\hat{\rho}]\coloneqq\textrm{Tr}\left[\hat{\rho}^{2}\right]$ being the purity of the state $\hat{\rho}$. The Eq. (44) above possesses a less trivial structure and dynamics, despite clearly approaching zero for $t\rightarrow\infty$, i.e., one can show that
$\psi^{2}(t)-4\left|Q(t)\right|^{2}=4\left(P_{0}-f\right)^{2}\frac{\left(1-e^{\gamma\left(t-\tau_{1/2}\right)}\right)^{2}}{e^{2\gamma t}},$ (45)
which is monotonic in time $t$ for $P_{0}\leq 1/2$ and non-monotonic otherwise. In the latter case, the minimum value is reached at the time $\tau_{1/2}$, satisfying $\psi^{2}(\tau_{1/2})-4\left|Q(\tau_{1/2})\right|^{2}=0$. This time dependency leads to a non-monotonic evolution of the coherent ergotropy $\mathcal{R}^{coh}[\hat{\rho}(t)]$, meaning that, despite interaction with the environment, the coherent ergotropy may initially increase before its inevitable decrease. This behavior has a direct impact on the battery’s discharging rate. Figure  12 illustrates the landscape of coherent ergotropy as a function of dissipation time and initial coherence $Q_{0}\in[0,\sqrt{P_{0}(1-P_{0})}]$ for (a) $P_{0}=0.2$ and (b) $P_{0}=0.8$. The dashed red lines indicate the curves $\tau_{1/2}$. As expected, in Fig. 12a, the coherent ergotropy decays monotonically over time. In contrast, in Fig. 12b, the previously described behavior is observed: $\mathcal{R}^{coh}[\hat{\rho}(t)]$ increases over $t\in[0,\tau_{1/2}]$, followed by a subsequent decay for $t>\tau_{1/2}$. It is also noteworthy that the contribution of the coherent component to the total ergotropy is smaller than that of the incoherent component.
Figure 12: Coherent ergotropy in terms of the dissipation time and initial coherence $Q_{0}\in[0,\sqrt{P_{0}(1-P_{0})}]$ for (a) $P_{0}=0.2$ and (b) $P_{0}=0.8$. The dashed red line highlights the curve $\tau_{1/2}$. It was assumed $\omega=1$, $f=0.3$, $\gamma=10^{-2}$.
@@ Appendix B Energy and ergotropy shifts
By design, the described protocol requires applying $\hat{M}_{m}$ and $\hat{W}_{w}$ to the battery state before and after some dissipation time. Both actions inherently cause energetic changes, resulting in additional charging or discharging of the battery. To ensure minimal external disturbance, it is essential that no net energy or ergotropy is imparted. In other words, the valid working regime of the protocol should be constrained to parameters that simultaneously satisfy $\varepsilon_{mw}(\tau)=0$ and $\mathcal{W}_{mw}(\tau)=0$. For fixed values of $\gamma$, $f$, $P_{0}$, $m$, and $\tau$, the former condition can be achieved by adjusting the reversal measurement strength to $\tilde{w}$, as defined in Eq. (24). Figure 13 illustrates the behavior of the $\tilde{w}$-surface for null energetic shift in terms of the initial excited population $P_{0}$ and strength $m$, evaluated at $\tau=\tau_{\gamma}$. The suppressed region represents non-physical values of measurement strength, where $\tilde{w}\notin[0,1]$. Higher values of $\tilde{w}$ are required as $m$ increases and for intermediate values of $P_{0}$. Since the post-weak measurement state remains invariant under the action of $\hat{M}_{m}$ for $P_{0}=0,1$ (implying $Q_{0}=0$), no energy is transferred during the initial measurement, so $\tilde{w}=0$ is sufficient to ensure $\varepsilon_{m\tilde{w}}(\tau)=0$. These general characteristics hold for different dissipation times $\tau$. Notably, for long times, the $\tilde{w}$-surface converges
@@
Figure 13: Null energetic shift surface in terms of $P_{0}$ and $m$, computed at $\tau_{\gamma}$. The non-physical region, $\tilde{w}\notin[0,1]$, is suppressed. It was assumed $f=0.3$ and $\gamma=10^{-2}$.
Nevertheless, given that $\varepsilon_{m\tilde{w}}(\tau)=0$, it remains necessary to identify parameters that also satisfy $\mathcal{W}_{m\tilde{w}}(\tau)=0$. To illustrate this, Fig. 14 presents the energetic shift profile across different parameter settings. Figures 14a–c display the behavior of the profile with respect to the reversal measurement strength $w$ and the dissipation time $\tau$, assuming an initial coherence $Q_{0}=Q_{max}\coloneqq\sqrt{P_{0}(1-P_{0})}$ and weak measurement strength $m=0.5$, for different initial population values: (a) $P_{0}=0.4$, (b) $P_{0}=0.8$, and (c) $P_{0}=0.9$. As expected, most of the parameter space shows a non-zero net energetic exchange. However, the intersection of the dashed black and continuous red curves marks the operational points of the protocol. The former indicates a null energetic shift, characterized by $\tilde{w}$, while the latter identifies points where $\mathcal{W}_{mw}(\tau)=0$ holds. For the chosen parameters, these intersections may only occur at specific dissipation times.
@@
Similarly, Figs. 14d–f present the energy shift landscape for the same initial populations, assuming $Q_{0}=0$ and equal measurement strengths, $m=w=\eta$. The dashed black lines mark points defined by the null energetic shift curves $\eta_{2,3}$, while the continuous red lines and hatched red areas highlight where a null ergotropy shift is obtained. Again, the intersections represent the desired working regime. As discussed in the main text, the $\eta_{2}$ curve for incoherent states leads to a cyclic transformation, $\hat{\rho}_{0}\rightarrow\hat{\tau}_{th}\rightarrow\hat{\rho}_{0}$, ensuring that the battery state remains unchanged during dissipation, thus satisfying $\varepsilon_{\eta_{2}\eta_{2}}=\mathcal{W}_{\eta_{2}\eta_{2}}=0$ for all $\tau$. However, for the selected parameters, the $\eta_{3}$ curve does not exhibit any operational points.
@@
Figure 14: Energy shift profile. (a)-(c) Energy shift in terms of the reversal strength $w$ and dissipation time $\tau$, assuming fixed strength $m=0.5$ and $Q_{0}=Q_{max}\coloneqq\sqrt{P_{0}(1-P_{0})}$, for (a) $P_{0}=0.4$, (b) $P_{0}=0.8$ and (c) $P_{0}=0.9$. The dashed black lines represent the points along $\tilde{w}$, given by Eq. (24), where $\varepsilon_{mw}(\tau)=0$, while the continuous red curves highlight where $\mathcal{W}_{mw}(\tau)=0$. (d)-(f) Energy shift for equal measurement strengths $m=w=\eta$ relative to $\tau$, assuming $Q_{0}=0$ for (d) $P_{0}=0.4$, (e) $P_{0}=0.8$ and (f) $P_{0}=0.9$. The black dashed lines represent the curves $\eta_{2}$ and $\eta_{3}$ for a null energetic shift, given by Eq. (25). The hatched red region and continuous red curves represent the points where $\mathcal{W}_{mw}(\tau)=0$ is satisfied. It was assumed $\omega=1$, $f=0.3$, $\gamma=10^{-2}$.

### v2

@@ VI Acknowledgements
The authors thank Michał Horodecki and Paweł Mazurek for the insightful discussions and suggestions, and the referees for the valuable inputs. A.H.A.M. and R.S. acknowledge support from National Science Centre, Poland Grant OPUS-21 (No. 2021/41/B/ST2/03207). B.A. acknowledges support from National Science Center, Poland within the QuantERA II Programme (No 2021/03/Y/ST2/00178, acronym ExTRaQT) that has received funding from the European Union’s Horizon 2020. P.R.D. acknowledges support from the NCN Poland, ChistEra-2023/05/Y/ST2/00005 under the project Modern Device Independent Cryptography (MoDIC).
@@ Appendix A Dissipation and battery discharging
Figure 17: Discharging process: Ergotropy behavior as a function of the dissipation time for distinct initial states, characterized by $P_{0}$ and $|Q_{0}|^{2}$. It was assumed $f=0.3$ and $\gamma=10^{-2}$.
@@
This phenomenon is illustrated in Fig. 17 for various coherent and incoherent initial states, characterized by $P_{0}$ and $|Q_{0}|^{2}$. Notably, when the initial state is active, the ergotropy decreases monotonically and irreversibly over time. In contrast, for passive states, the ergotropy remains constant and zero, as indicated by the red dashed line. Additionally, note that the discharging rates vary significantly. For instance, in the case of a fully charged battery, represented by the solid black line, the decay is faster than for the initially coherent states. From this perspective, certain partially charged initial states may retain more ergotropy after a period of dissipation, presenting a potential advantage. This counterintuitive behavior arises from the complex interplay between the incoherent and coherent components of ergotropy during thermalization, resembling the ergotropic counterpart of the Mpemba effect [140, 141].
Figure 18: Incoherent ergotropy in terms of the dissipation time and initial population $P_{0}$. The dashed red line highlights the curve $\tau_{1/2}$, representing the instant where the diagonal state becomes passive ($P(\tau_{1/2})=1/2$). It was assumed $f=0.3$, $\gamma=10^{-2}$.
@@
On the other hand, the coherent ergotropy at time $t$ is written according to Eq. (15). As elucidated in the main text, it possesses a less trivial structure and dynamics, which is monotonic for $P_{0}\leq 1/2$ and non-monotonic otherwise. Thus, despite environmental interaction, $\mathcal{R}^{coh}[\hat{\rho}(t)]$ may initially increase before eventually decreasing for $t>\tau_{1/2}$, which significantly influence the battery’s discharging rate. To better access the dynamics of the coherent ergotropy, Figure 19 illustrates its landscape as a function of dissipation time and initial coherence $|Q_{0}|^{2}\in[0,P_{0}(1-P_{0})]$ for (a) $P_{0}=0.2$ and (b) $P_{0}=0.8$. The dashed red lines indicate the curves $\tau_{1/2}$.
Figure 19: Coherent ergotropy in terms of the dissipation time and initial coherence $|Q_{0}|^{2}\in[0,P_{0}(1-P_{0})]$ for (a) $P_{0}=0.2$ and (b) $P_{0}=0.8$. The dashed red line highlights the curve $\tau_{1/2}$. It was assumed $f=0.3$, $\gamma=10^{-2}$.
As expected, in Fig. 19a, the coherent ergotropy decays monotonically over time while $P(t)$ approaches $f$. In contrast, in Fig. 19b, the previously described behavior is observed: $\mathcal{R}^{coh}[\hat{\rho}(t)]$ increases over $t\in[0,\tau_{1/2}]$, followed by a subsequent decay for $t>\tau_{1/2}$. It is also noteworthy that the contribution of the coherent component to the total ergotropy is smaller than that of the incoherent component.
@@ Appendix B Energy and ergotropy shifts
By design, the described protocol requires the application of $\hat{M}_{m}$ and $\hat{W}_{w}$ to the battery state before and after some dissipation time. Both actions inherently cause energetic changes, resulting in additional charging or discharging of the battery. To ensure minimal external disturbance, it is essential that no net energy or ergotropy is imparted. In other words, the valid working regime of the protocol should be constrained to parameters that simultaneously satisfy $\varepsilon_{mw}(\tau)=0$ and $\mathcal{W}_{mw}(\tau)=-\varepsilon_{mw}^{p}(\tau)=0$. For fixed values of $\gamma$, $f$, $P_{0}$, $m$, and $\tau$, the former condition can be achieved by adjusting the reversal measurement strength to $\tilde{w}$, as defined in Eq. (24). Figure 20 illustrates the behavior of the $\tilde{w}$-surface for a null energetic shift in terms of the initial excited population $P_{0}$ and strength $m$, evaluated at $\tau=\tau_{\gamma}$ for different temperatures (encoded in $f$) given by (a) $f=0$ and (b) $f=0.30$. The suppressed region in (b) represents non-physical values of measurement strength, where $\tilde{w}\notin[0,1]$. In general, higher values of $\tilde{w}$ are required as $m$ increases. However, its behavior in terms of $P_{0}$ clearly depends on the temperature, especially for $P_{0}<0.5$. In this sense, for $f=0.30$ the stronger values of $\tilde{w}$ are constrained to the intermediate values of $P_{0}$, while this region increases for lower temperatures toward small $P_{0}$. Also, for both cases, one should have $\tilde{w}=0$ to ensure that $\varepsilon_{m\tilde{w}}(\tau)=0$ for $P_{0}=1$ and $P_{0}=0$ (not visible in (a)), i.e., since the post-weak measurement state remains invariant under the action of $\hat{M}_{m}$ no energy is transferred during the initial measurement, which means the reversal measurement strength should be null. These general characteristics hold for different dissipation times $\tau$. Notably, for long times and $f\neq 0$, the $\tilde{w}$-surface converges
@@
For $f=0$ and $\tau\rightarrow\infty$, the battery thermalizes to the ground state, such that the reversal measurement has no effect and the only energetic contribution is due to the first weak measurement. In this case, the only way to guarantee a null energetic shift is by $m=0$ or $P_{0}=0,1$.
@@
Figure 20: Null energetic shift $\tilde{w}$-surface in terms of $P_{0}$ and $m$, computed at $\tau_{\gamma}$ for (a) $f=0$ and (b) $f=0.3$. The non-physical region, $\tilde{w}\notin[0,1]$, is suppressed. It was assumed $\gamma=10^{-2}$.
Nevertheless, since $\varepsilon_{m\tilde{w}}(\tau)=0$, it is crucial to identify the parameters that also satisfy $\mathcal{W}_{m\tilde{w}}(\tau)=-\varepsilon_{m\tilde{w}}^{p}(\tau)=0$. To illustrate this and the influence of temperature, Fig. (21) presents the ergotropy shift profile at $\tau_{\gamma}$ for different parameter settings. Figures (21)a–c show how the ergotropy shift depends on the initial population $P_{0}$ and measurement strength $m$, assuming maximal initial coherence $|Q_{0}|^{2}=|Q_{\max}|^{2}=P_{0}(1-P_{0})$. These plots correspond to different temperatures: (a) $f=0$, (b) $f=0.15$, and (c) $f=0.30$. The desired condition $\mathcal{W}_{m\tilde{w}}(\tau)=0$ (marked by red curves) is not satisfied for any $P_{0}$ and $m$ when $f=0$, indicating that at zero temperature, the protocol inevitably causes nonzero ergotropy shift. However, as $f$ increases, the number of operational points grows, suggesting that thermal effects help to balance the ergotropy changes. Additionally, the black and blue dashed curves indicate the null energy shift curves $\eta_{2,3}$, revealing the interplay between ergotropy and energy changes. Similarly, Figs. 21d–f illustrate the ergotropy shift for the same temperatures but with $|Q_{0}|^{2}=0$, revealing a significantly different behavior. As the temperature increases, the regions satisfying $\mathcal{W}_{m\tilde{w}}(\tau)=0$ (red hatched regions) expand, particularly for larger $P_{0}$. This suggests that higher temperatures introduce a stabilizing effect, enabling the protocol to operate within the required conditions over a broader parameter range, reinforcing the idea that thermal fluctuations can counteract undesired energy and ergotropy shifts.
Figure 21: Ergotropy shift profile at time $\tau_{\gamma}$ along $\tilde{w}$. (a)-(c) Ergotropy shift in terms of the initial population $P_{0}$ and measurement strength $m$, assuming $|Q_{0}|^{2}=|Q_{max}|^{2}=P_{0}(1-P_{0})$ for each $P_{0}$, for (a) $f=0$, (b) $f=0.15$ and (c) $f=0.3$. (d)-(f) Ergotropy shift in terms of the initial population $P_{0}$ and measurement strength $m$, assuming $|Q_{0}|^{2}=0$ for (d) $f=0$, (e) $f=0.15$ and (f) $f=0.3$. The hatched red region and continuous red curves highlight where $\mathcal{W}_{m\tilde{w}}(\tau)=0$. The black and blue dashed curves represent contours of the energy shift $\eta$, revealing the interplay between ergotropy and energy transfer. It was assumed $\gamma=10^{-2}$.
@@
As mentioned above, the operational points of the TWM protocol are defined by the simultaneous fulfillment of both null energetic and ergotropic shifts. In this sense, Fig. 22 presents the energetic shift profile for different parameters. Figures 22a–c display the behavior of the profile with respect to the reversal measurement strength $w$ and the dissipation time $\tau$, assuming $f=0.30$, an initial coherence $|Q_{0}|^{2}=|Q_{max}|^{2}=P_{0}(1-P_{0})$, and weak measurement strength $m=0.5$, for different initial population values: (a) $P_{0}=0.4$, (b) $P_{0}=0.8$, and (c) $P_{0}=0.9$. As expected, most of the parameter space shows a non-zero net energetic exchange. However, the intersection of the dashed black and continuous red curves marks the operational points of the protocol. The former indicates a null energetic shift, characterized by $\tilde{w}$, while the latter identifies points where $\mathcal{W}_{mw}(\tau)=0$ holds. For the chosen parameters, these intersections may only occur at specific dissipation times. Analogously, Figs. 22d–f present the energy shift landscape for the same initial populations, assuming $|Q_{0}|^{2}=0$ and equal measurement strengths, $m=w=\eta$. The dashed black lines mark points defined by the null energetic shift curves $\eta_{2,3}$, while the continuous red lines and hatched red areas highlight where a null ergotropy shift is obtained. Again, the intersections represent the desired working regime. As discussed in the main text, the $\eta_{2}$ curve for incoherent states leads to a cyclic transformation, $\hat{\rho}_{0}\rightarrow\hat{\tau}_{th}\rightarrow\hat{\rho}_{0}$, ensuring that the battery state remains unchanged during dissipation, thus satisfying $\varepsilon_{\eta_{2}\eta_{2}}=\mathcal{W}_{\eta_{2}\eta_{2}}=0$ for all $\tau$. However, for the selected parameters, the $\eta_{3}$ curve does not exhibit any operational points.
@@
Figure 22: Energy shift profile. (a)-(c) Energy shift in terms of the reversal strength $w$ and dissipation time $\tau$, assuming fixed strength $m=0.5$ and $|Q_{0}|^{2}=|Q_{max}|^{2}=P_{0}(1-P_{0})$, for (a) $P_{0}=0.4$, (b) $P_{0}=0.8$ and (c) $P_{0}=0.9$. The dashed black lines represent the points along $\tilde{w}$, given by Eq. (24), where $\varepsilon_{mw}(\tau)=0$, while the continuous red curves highlight where $\mathcal{W}_{mw}(\tau)=0$. (d)-(f) Energy shift for equal measurement strengths $m=w=\eta$ relative to $\tau$, assuming $|Q_{0}|^{2}=0$ for (d) $P_{0}=0.4$, (e) $P_{0}=0.8$ and (f) $P_{0}=0.9$. The black dashed lines represent the curves $\eta_{2}$ and $\eta_{3}$ for a null energetic shift, given by Eq. (25). The hatched red region and continuous red curves represent the points where $\mathcal{W}_{mw}(\tau)=0$ is satisfied. It was assumed $f=0.3$, $\gamma=10^{-2}$.
Appendix C Two-qubit open quantum battery dynamics: The $X$-state
Figure 23: Concurrence dynamics for (a) $q=0.1$, $m_{1}=0.5$, $m_{2}\approx 0.6$, $w_{1}=w_{2}\approx 0.21$. Changes with (dashed blue) and without (continuous red) the protocol. The result demonstrates that weak measurements can enhance the entanglement generated by dissipation. (b) $q=0.9$, $m_{1}=0.5$, $m_{2}=0.9$, $w_{1}\approx 0.97$, $w_{2}=0.17$. Changes with (dashed blue) and without (continuous red) the protocol. The result shows that weak measurements while preserving ergotropy in this specific state, diminish entanglement when it is already present before dissipation takes place.
[…]
$\hat{\rho}_{X}(q)=\left(\begin{array}[]{cccc}q/2&0&0&q^{2}-q/2\\ 0&q(1-q)&0&0\\ 0&0&(1-q)^{2}&0\\ q^{2}-q/2&0&0&q/2\end{array}\right).$ (50)
This particular unitary operation was shown to maximize the quantum correlations presented in the final state. Moreover, entanglement (measured using concurrence, relative entropy, or entanglement negativity) vanishes at a finite critical temperature, other notions of quantum correlations—such as global quantum discord, local quantum uncertainty [159], and realism-based nonlocality [160]—gradually decay as the temperature increases.
As discussed in the main text, the dynamics of a two-qubit quantum battery in contact with a common environment is given by the following master equation
where the Hamiltonian reads $\hat{H}=\omega|e\rangle\langle e|_{1}+\omega|e\rangle\langle e|_{2}+J\left(\sigma_{+}^{1}\sigma_{-}^{2}+\sigma_{-}^{1}\sigma_{+}^{2}\right)$, with $J$ being the qubit’s coupling strength, and the dissipators written as
$\begin{split}\mathcal{D}_{\downarrow}[\bm{\cdot}]&\coloneqq\gamma(1-f)\left(\hat{S}_{+}\bm{\cdot}\hat{S}_{-}-\frac{1}{2}\left\{\hat{S}_{-}\hat{S}_{+},\bm{\cdot}\right\}\right),\\ \mathcal{D}_{\uparrow}[\bm{\cdot}]&\coloneqq\gamma f\left(\hat{S}_{-}\bm{\cdot}\hat{S}_{+}-\frac{1}{2}\left\{\hat{S}_{+}\hat{S}_{-},\bm{\cdot}\right\}\right),\end{split}$ (52)
where $\hat{S}_{\pm}=\hat{\sigma}_{\pm}^{1}+\hat{\sigma}_{\pm}^{2}$ are the raising and lowering ladder operators. Assuming as initial state $\hat{\rho}_{0}$, the $X$-state from Eq. (50) one can show that for all $t\geq 0$, the density matrix is written as
$\hat{\rho}(t)=\left(\begin{array}[]{cccc}\rho_{11}(t)&0&0&\rho_{14}(t)\\ 0&\rho_{22}(t)&\rho_{23}(t)&0\\ 0&\rho_{23}^{*}(t)&\rho_{33}(t)&0\\ \rho_{14}^{*}(t)&0&0&\rho_{44}(t)\end{array}\right),$ (53)
where, according to Eq. (51), the dynamics is given by the following set of coupled differential equations:
$\begin{split}\frac{d\rho_{11}(t)}{dt}&=-2\gamma f\rho_{11}(t)-\gamma(f-1)\left[\rho_{22}(t)+2\textrm{Re}\left[\rho_{23}(t)\right]+\rho_{33}(t)\right],\\ \frac{d\rho_{14}(t)}{dt}&=-\left(\gamma-2i\omega\right)\rho_{14}(t),\\ \frac{d\rho_{22}(t)}{dt}&=\gamma f\rho_{11}(t)-\gamma\left[\rho_{22}(t)+\left(f-1\right)\rho_{44}(t)\right]-\left(\gamma+2iJ\right)\textrm{Re}\left[\rho_{23}(t)\right]+2iJ\rho_{23}(t),\\ \frac{d\rho_{23}(t)}{dt}&=\frac{1}{2}\gamma(2f\rho_{11}(t)-\rho_{22}(t)-\rho_{33}(t)-2(f-1)\rho_{44}(t)-2\rho_{23}(t))+iJ(\rho_{22}(t)-\rho_{33}(t)),\\ \frac{d\rho_{33}(t)}{dt}&=\gamma f\rho_{11}(t)-\gamma\left[\rho_{33}(t)+\left(f-1\right)\rho_{44}(t)\right]-\left(\gamma-2iJ\right)\textrm{Re}\left[\rho_{23}(t)\right]-2iJ\rho_{23}(t),\\ \frac{d\rho_{44}(t)}{dt}&=2\gamma\left(f-1\right)\rho_{44}(t)+\gamma f\left(\rho_{22}(t)+2\textrm{Re}\left[\rho_{23}(t)\right]+\rho_{33}(t)\right).\end{split}$ (54)
Note that the non-diagonal elements $\rho_{23}(t)$ and $\rho_{23}^{*}(t)$ are, in general, non-null during the dissipative dynamics. More importantly, the solution of previous equations allows one to compute the time-evolution of the relevant quantities in terms of $q$, in particular ergotropy, as presented in Section IV.1. In this sense, the concurrence change along the protocol execution for the examples without ($q=0.1$) and with ($q=0.9$) initial entanglement is presented in Figs. 23 (a) and (b), respectively. For the former, the dynamic-induced entanglement observed for the purely dissipative time evolution is shown to be enhanced with the execution of the TWM protocol. For the latter, the entanglement is shown to be consumed faster compared to the case without the weak measurements. Nevertheless, in both scenarios, the use of the TWM protocol mitigates ergotropy loss.